OpenStudy (anonymous):

Can anyone help Me figure out how I'm supposed to check these? I'll fan and give medal

OpenStudy (anonymous):

Here's an example in my book but I still don't understand

OpenStudy (campbell_st):

so what has happened in the example is that initially the equation -x + 3 = -2x + 5 so you need to solve for x the the checking part is using the solution to see if it makes sense this is done by replacing x in the original equation with 2 and seeing if it makes sense. hope that helps

OpenStudy (campbell_st):

so if you have the equation -x + 3 = -2x + 5 add 2x to both sides 2x 2x ------------------ x + 3 = 5 now subtract 3 from both sides -3 -3 ------------ x = 2 so that's the solution. Put it into the original equation -2 + 3 = -2x2 + 5 do the arithmetic 1 = 1 so the solution x = 2 is correct...

OpenStudy (anonymous):

@campbell_st I'm so confused can you at least walk me through the first problem so I will understand what I've wasted 30 minutes on(I waste time a lot)

OpenStudy (campbell_st):

ok... so the equation is -x + 3 = -2x + 5 and you need to solve for x... any thoughts on the 1st step..?

OpenStudy (campbell_st):

or just post the 1st problem

OpenStudy (anonymous):

Oh i just realizedthe pic i posted of my questions didnt go through @campbell_st

OpenStudy (campbell_st):

ok

OpenStudy (campbell_st):

can you repost the question

OpenStudy (anonymous):

yeah i was trying to upload the photo but its not working so I'm just gonna use the site

OpenStudy (anonymous):

\[3x + 7 = 16 + 2x\]

OpenStudy (campbell_st):

ok... so any thoughts...?

OpenStudy (anonymous):

My work for solving them was 1) 3x - 2x = 16 - 7 2) x= 16 - 7 3) x= 9

OpenStudy (anonymous):

I solved the equation by myself before I came back on here but its the "Check" part that I'm confused about

OpenStudy (campbell_st):

ok.. so the solution is perfect... the check idea is to simply substitute your solution, x = 9 back into the original equation to see if its true \[does~~~3 \times 9 + 7 =16 + 2 \times 9\] if it does then you know your solution is right

OpenStudy (anonymous):

@campbell_st thank you so much this makes way more sense now